The Vapour Density of the equilibrium mixture of the reaction : SO2Cl2(g) --> SO2(g) + Cl2(g)is 50. The % dissociation of SO2Cl2 is?

Grade:12

3 Answers

Vikas TU
7 years ago
Already Answered this:
Check your previous qstn.
ANswer=>
Degree of dissociation from the formulae:
1 + a = Mth./Mobs.
Mobs. = 2*V.D = > 100
hecne,
a = 135/100 – 1 => 0.35
percentage of dissociation = > 35% would be the degree of dissociation.
Aswanth
5 years ago
Degree of dissociation from the formulae:
1 + a = Mth./Mobs.
Mobs. = 2*V.D = > 100
hecne,
a = 135/100 – 1 => 0.35
percentage of dissociation = > 35% would be the degree of dissociation.
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Degree of dissociation from the formulae:
1 + a = Mtheoritical / Mobserbed
Mobserved = 2*V.D = 100
hence, a = 135/100 – 1 = 0.35
percentage of dissociation = 35% would be the degree of dissociation

Thanks and Regards